Meiger Mining, Inc., has just discovered two new mining sites for iron ore. Geologists and engineers have come up with the estimates on the following page regarding costs and ore yields if the mines are opened.

Site A Site B
Variable extraction costs per ton $ 3.80 $ 4.00
Fixed costs over the life of the mine:
Blasting $ 150,000 $ 185,000
Construction 225,000 240,000
Maintenance 25,000 20,000
Restoration costs 40,000 35,000
Total fixed costs $ 440,000 $ 480,000
Total tons of ore that can be extracted over the life of the mine: 200,000 160,000

Meiger’s owners currently demand a return of 20 percent of the market price of iron ore.

Required:

a. If the current market price of iron ore is $8 per ton, what is Meiger’s target cost per ton?

b-1. Given the $8 market price, compute the total cost per ton for Site A and Site B if fixed overhead costs are assigned to products on the basis of estimated tons of ore that can be extracted over the life of the mine.

b-2. Should either of the mines be opened?

c-1. The engineer working on Site B believes that if a custom conveyor system is installed, the variable extraction cost could be reduced to $3 per ton. The purchase price of the system is $25,000, but the costs to restore the site will increase to $45,000 if it is installed. Given the current $8 market price, compute the total cost per ton for site B.

c-2. Should Meiger install the conveyor and open Site B?
Answer #1

a. If the current market price of iron ore is $8 per ton, then Meiger's target cost per ton is,

$8.0 - ($8.0X 20%) = $6.4 per Ton

b-1. Total cost per ton:

For site A :

Variable cost per ton +Fixed cost per ton

$3.80 +($440,000÷ 200,000 tons) = $6 per ton

For Site B:

$4.00 +( $480,000 ÷ 160,000) = $7 per ton

b-2: to meet the demand of 20% only site A shoudl be opened because its cost per ton is less then the targeted cosst of $6.40 and since the cost per ton for site B is $7 it will not give the desired return.

c. Total Cost per ton for site B:

Variable cost per ton +Fixed cost per ton

$3.00 +[$185,000+240,000+20,000+25,000+45,000)÷ 160,000tons] = $6.22 per ton

Current Market Price: $8.00 , so target cost per ton is,

$8.00 - ($8.00X 20%) = $6.4 per Ton

Yes, Meiger should install the conveyor and open Site B,because the cost per ton after installing of conveyer is less then the targeted cost of $6.40.